OpenStudy (anonymous):

Solve the differential Eqn: y'=(x+y-2)Whole square

OpenStudy (unklerhaukus):

\[y'=(x-y-2)^2\]

Parth (parthkohli):

\[y' = (x + y - 2)^2 \]

OpenStudy (unklerhaukus):

dammit

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i'm lookin for solution...

OpenStudy (unklerhaukus):

it dosent look separable or linear

OpenStudy (unklerhaukus):

it dosent look homogenous

OpenStudy (anonymous):

yeah u'r right

OpenStudy (unklerhaukus):

is it a Bernoulli equation?

OpenStudy (unklerhaukus):

if we can write the equation in this form \[y' + p(x)y = q(x)y^n\]it is a bernoulli equation

OpenStudy (anonymous):

I don't think it is...

OpenStudy (unklerhaukus):

i guess we have to make a substitution

OpenStudy (unklerhaukus):

let \(w=x+y(x)\) \(\text dw=\)

OpenStudy (anonymous):

yep you got it

OpenStudy (anonymous):

@UnkleRhaukus still there?

OpenStudy (anonymous):

Hah that was a trip down memory lane pretty fun:)

OpenStudy (anonymous):

@shubham9495 still there?

OpenStudy (unklerhaukus):

hmm..

OpenStudy (anonymous):

your sub.s works

OpenStudy (anonymous):

w= x+y w' = y' +1 y' = w' -1 w'-1 = (w-2)^2 w'-1 = w^2 -4w +4 w'= w^2 -4w +5 dw/(w^2 -4w +5) = dx

OpenStudy (anonymous):

integrate -arctan(2-w) = x 2-w = tan(-x) w= -tan(-x) +2 x+y = -tan(-x) +2 y= -tan(-x) {{call this tan(x) now}} +2 -x

OpenStudy (anonymous):

verify in original tan^2(x) = tan^2(x)

OpenStudy (unklerhaukus):

what happened to the constant ?

OpenStudy (anonymous):

I was thinking about that...

OpenStudy (anonymous):

think it has to be a whole period , so it's just a phase shift... not %100 on that however

OpenStudy (unklerhaukus):

the solution to a differential equation should have a constant

OpenStudy (anonymous):

I was talking abut the constant of integration

OpenStudy (anonymous):

Guess I should have left it in... rather than assume it was a whole period shift